### v2

@@ 4 Consistent conjectures
In this section, we present a sufficient condition on the family of conjectures, $\Phi$, such that the set of $\Phi$-solutions is nonempty (Theorem 4.1). We dub the conjectures satisfying the sufficient condition as consistent conjectures. Consistent conjectures guarantee that a candidate matching, defined below, is a solution. In what follows, we first introduce the candidate matching(s), and then we state Theorem 4.1. After discussing the intuition behind the result, we connect it to similar observations in Hafalir (2008) and Doval (2022). We conclude this section by showing that the extension to dynamic matching markets of a standard solution concept in matching with externalities, rational expectations, does not entail consistent conjectures.
@@
    Preferences $\{u_{\varphi_{1}}(a,\cdot):B_{1}\cup\{a\}\mapsto\mathbb{R}:a\in A_{1}\},\{v_{\varphi_{1}}(\cdot,b):A_{1}\cup\{b\}\mapsto\mathbb{R}:b\in B_{1}\}$ are given by:
@@
      For each agent $a\in A_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,a)$ only if $u(a,b)\geq\min\{U_{1}(a,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,a)\}$. Moreover, for $b,b^{\prime}\in B_{1}$, $u_{\varphi_{1}}(a,b)\geq u_{\varphi_{1}}(a,b^{\prime})$ only if $u(a,b)\geq u(a,b^{\prime})$.
@@
      For each agent $b\in B_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(b,b)$ only if $v(a,b)\geq\min\{V_{1}(b,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,b)\}$. Moreover, for $a,a^{\prime}\in A_{1}$, $v_{\varphi_{1}}(a,b)\geq v_{\varphi_{1}}(a^{\prime},b)$ only if $v(a,b)\geq v(a^{\prime},b)$.
@@
In the one-period economy induced by the conjectures $\varphi_{1}$, agents’ preferences are truncated using their worst conjectured matching. That is, in the one-period economy induced by $\varphi_{1}$, an agent $a$ ranks matching partners on side $B$ according to $u(a,\cdot)$, except that agents on side $B$ who are worse than the worst conjectured continuation matching are now deemed unacceptable. In other words, in the one-period economy induced by $\varphi_{1}$, each agent in $A_{1}\cup B_{1}$ has as an outside option the worst conjectured matching.
@@
Suppose we know that $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence and consider the following set of candidate matchings:
@@
In words, a candidate matching, $m^{\star}\in M^{\star}$ satisfies the following. In period $1$, $m_{1}^{\star}$ is a stable matching for $E_{1,\varphi_{1}}$, i.e., $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$. Because the one-period economy induced by $\varphi_{1}$ is devoid of any externalities, its set of stable matchings is nonempty and hence, at least one such matching exists (Gale and Shapley, 1962). Given the period-1 matching, $m^{\star}_{1}$, the continuation matching, $(m^{\star}_{t})_{t=2}^{T}$ is selected from $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))$, which is nonempty by assumption. Because $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$ may not be a singleton and given one such stable matching, multiple continuation matchings may exist in the solution set of the induced economy, many candidates matchings may exist.Footnote 8Footnote 8Footnote 8Whereas many stable matchings for the one-period economy induced by the period-1 conjectures may exist, the Lone Wolf theorem (McVitie and Wilson, 1970) implies they all have the same set of unmatched agents and hence, they all induce the same continuation economy.
Any matching in $M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$ is a natural candidate to be a $\Phi$-solution because it satisfies most of the conditions in 7. To see this, fix any such matching, $m^{\star}$, and note the following. First, $m^{\star}$ satisfies part 4 of 7 by construction. Second, $m^{\star}$ satisfies parts 1, 2, and 3 of 7 for all period-1 agents $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)\neq k$. By definition of $E_{1,\varphi_{1}}$, every agent who matches in $t=1$ is matched to someone better than their worst conjectured matching. Furthermore, because $m^{\star}_{1}$ is chosen from $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, no pair of matched agents under $m^{\star}_{1}$ can form a blocking pair.
@@
As we explain next, the only reason any such matching $m^{\star}$ may fail to be a $\Phi$-solution is that a “gap” exists between the conjectures of an agent who is unmatched in $t=1$ under $m^{\star}$ and the matching $m^{\star}$. To see this, note that the above discussion implies that if $m^{\star}$ is not a $\Phi$-solution, then $m^{\star}$ has a block involving a period-$1$ agent, $k$, who is unmatched in $t=1$, i.e., $m^{\star}_{1}(k)=k$. We claim next that if $k$ blocks $m^{\star}$, then $m^{\star}$ is not a conjecture for $k$. To fix ideas, let $k=a\in A_{1}$ and suppose $a$ blocks $m^{\star}_{1}$ with an agent $b\in B_{1}$. Because $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, it must be that at least one of $a$ or $b$ prefer their worst conjectured matching to matching with each other; assume for simplicity that this holds for $a$. Because $a$ wants to block $m^{\star}$ with $b$, it must be that $a$’s outcome under $m^{\star}$ is worse than the worst matching in $\varphi_{1}(\emptyset,a)$. In other words, when considering what matchings may ensue in the continuation economy, $a$ does not deem $m^{\star}$ a feasible conjecture. Because $a$ only cares about the continuation matching $(m^{\star}_{t})_{t=2}^{T}$, that $m^{\star}$ is not a feasible conjecture is akin to agent $a$’s conjectures not being consistent with the solution concept in the continuation. If instead $a$’s conjectures were consistent with the solution concept in the continuation, then they would not be willing to block $m^{\star}$ with $b$.
The above discussion motivates the result in Theorem 4.1:
@@
Suppose $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence. Fix $E^{T}$ and a candidate matching $m^{\star}\in M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$. Suppose that for all $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$, we have that
$m^{\star}\in\varphi_{1}(\emptyset,k).$ (CC)
Then, $m^{\star}\in\Sigma_{T}^{\Phi}(E^{T})$.
The proof of Theorem 4.1 and other results can be found in the appendix.
In words, if for some candidate matching $m^{\star}$, $m^{\star}$ is a conjectured matching for each of the unmatched agents under $m^{\star}$ in period $1$, then $m^{\star}$ is a solution for the economy. Whereas agents who match in period $1$ under $m^{\star}$ may have different conjectures about the matching that would ensue conditional on remaining unmatched, consistency requires that enough commonality exists in the conjectures of those agents who remain unmatched. Indeed, all agents who remain unmatched in period $1$ should agree that $(m^{\star}_{t})_{t=2}^{T}$ is a feasible continuation matching.
@@
As stated, the condition in Theorem 4.1 seems difficult to check without actually computing $m^{\star}$. Yet, Theorem 4.1 suggests a family of conjectures that always lead to nonempty solutions without actually computing $m^{\star}$. Indeed, it follows from Theorem 4.1 that if each period-1 agent conjectures that from $t=2$ onward, the continuation matching is an element of $\Sigma_{T-1}^{\Phi}(\cdot)$, then $\Sigma_{T}^{\Phi}$ is nonempty valued. Formally, define for each $k\in A_{1}\cup B_{1}$,
$\varphi^{*}_{1}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(\overline{m}_{1}))\right\}.$ (4.1)
@@
These conjectures reflect that $k$ believes that the economy abides by the set of $\Phi$-solutions from period $2$ onward. Thus, for each conjectured period-1 matching, $\overline{m}_{1}$, $k$ anticipates that the remaining unmatched agents and the new entrants will object to any matching $(\overline{m}_{t})_{t=2}^{T}$ that does not conform with $\Sigma_{T-1}^{\Phi}(\overline{m}_{1})$.Footnote 9Footnote 9Footnote 9These conjectures also appear in the solution concept for two-sided, repeated matching markets in Kotowski (2019). For instance, if $T=2$, these conjectures rule out matchings $\overline{m}$ such that the period-2 matching, $\overline{m}_{2}$ is not stable in the induced economy, $E_{2}^{2}(\overline{m}_{1})$. Interestingly, that these conjectures induce a nonempty solution concept implies that the perfection requirement alone is compatible with the existence of solutions for dynamic matching markets. Instead, difficulties may arise if the conjectures also place structure on the period-$t$ matching, as we do in Section 5.
@@
Finally, even if the statement of Theorem 4.1 relies on the recursive definition of the set of $\Phi$-solutions for a length-$T$ economy taking the non-emptiness of $\Sigma_{T-1}^{\Phi}$ as given, the procedure to construct $m^{\star}$ and the statement of Theorem 4.1 can be expressed solely in terms of the model primitives. We do this in Appendix A for completeness. Here we note that this only leads to requiring consistency of the conjectures in each period along the path of $m^{\star}$: Given $m^{\star}$, consistency is defined as the requirement that for all $t\geq 1$, all $k\in\pazocal{A}_{t}(m^{\star^{t-1}})\cup\pazocal{B}_{t}(m^{\star^{t-1}})$ such that $m^{\star}_{t}(k)=k$, then $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k)$ (see Theorem A.1). Similarly, we recursively build the conjectures and the solution concept corresponding to Equation 4.1 and show that a solution concept built off of conjectures that rule out matchings that do not conform with the solution concept in the continuation economy is always nonempty (see A.1 in Appendix A).
@@
Connection with Hafalir (2008) and Doval (2022)
Hafalir (2008) studies stability in two-sided, one-to-one, matching markets with payoff externalities. The starting point of his analysis is that the negative results in Sasaki and Toda (1996) follow from conjectures being exogenously defined without regard to the economy at hand. The first main result in Hafalir (2008) provides a sufficient condition on the conjectures for a stable matching to exist in his environment. To explain the condition, we review the construction in Hafalir (2008). Using the conjectures, Hafalir (2008) constructs a two-sided economy without externalities as in Sasaki and Toda (1996). In present notation, consider a one-period economy with agents in $A_{1}\cup B_{1}$. For each pair of agents, $(a,b)\in A_{1}\times B_{1}$, a conjecture $\varphi(a,b)$ is a subset of the set of matchings such that $(a,b)$ is matched.Footnote 10Footnote 10Footnote 10Hafalir (2008) does not consider blocks by individual agents. Given the conjectures, say $a$ prefers $b$ over $b^{\prime}$ if $a$ prefers the worst matching in $\varphi$$(a,b)$ over the worst matching in $\varphi(a,b^{\prime})$. The preferences for agents on side $B$ are similarly defined. Under these induced preferences, a stable matching is guaranteed to exist. Fix any such matching, $\overline{m}$. Hafalir (2008, Proposition 2) shows that if for all pairs $(a,b)$ such that $\overline{m}(a)=b$, we have that $\overline{m}\in\varphi(a,b)$, then $\overline{m}$ is stable in the economy with externalities.
@@
Theorem 4.1 identifies a key property that underlies the existence result in Doval (2022). In fact, Remark 2 in that paper points out that the conjectures corresponding to dynamic stability satisfy condition CC for any dynamically stable matching–not just the corresponding candidate matchings $m^{\star}$–and discusses the key role this property plays in the existence result. In Doval (2022), the conjectures are defined recursively based on what the solution concept specifies in the continuation, similar to the conjectures in Equation 4.1. Because our framework isolates the conjectures from the solution concept, it allows us to identify the role of consistency without reference as to how the conjectures relate to the solution concept. In doing so, the result offers an alternative way to construct solution concepts for dynamic matching markets by focusing attention on guaranteeing the conjectures are consistent. Section 5 illustrates this approach.
@@
First, it is instructive to consider the implications of not having consistent conjectures in the context of the candidate matchings. To fix ideas, consider one such $m^{\star}$ and suppose an unmatched agent on side $A$, $a$, wishes to block $m^{\star}$ with an agent on side $B$, $b$. Suppose also that $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$, and prefers matching with agent $b$ to $m^{\star}$, so that $m^{\star}\notin\varphi_{1}(\emptyset,a)$. There is a sense in which $a$’s block is not credible: Whereas $a$ prefers to match with $b$ over $m^{\star}$, as soon as $a$ matches with $b$, $a$ would prefer to break up with $b$. After all, $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$. Thus, whenever the failure of consistency rules out a candidate matching $m^{\star}$ as a solution, this may be because the solution concept allows for these non-credible blocks.
To make the second point, considering a more general definition of consistency which dispenses with the candidate matchings is useful. Given a family of conjectures, $\Phi$, and the induced solution concept $\{\Sigma_{t}^{\Phi}:t\in\{1,\dots,T\}\}$, we say the solution concept satisfies generalized consistency if the following holds. For all economies $E^{T}$ and $\Phi$-solutions $m$ for $E^{T}$, for all periods $t$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, if $m_{t}(k)=k$, then $m\in\varphi_{t}(m^{t-1},k)$. That is, we require that condition CC holds not necessarily for the candidate matching $m^{\star}$, which may or may not be a solution, but for the actual solutions (provided they exist).
Suppose we know the set of $\Phi$-solutions is nonempty, but does not satisfy generalized consistency. For concreteness, let $m$ be a solution for $E^{T}$ for which a period $t\geq 1$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $m_{t}(k)\neq k$ and $m\notin\varphi_{t}(m^{t-1},k)$. Because $m$ is a solution, the issue of credible blocks discussed above does not arise. However, that $k$ does not conceive of $m$ as a potential outcome for the continuation economy suggests a tension exists between what the solution concept prescribes in the continuation economy in which $k$ is one of the agents available to match, and what $k$ expects would happen if they would object to $m$. That $m$ is not part of $k$’s conjectures means $k$ believes that $m$ cannot be enforced in the continuation. Thus, if the solution concept captures the set of self-enforcing outcomes in the continuation economy and $k$ agrees with this assessment, then $m$ should be part of $k$’s conjectures and the solution concept should satisfy generalized consistency. As we illustrate below, the rational expectations solution fails both consistency and generalized consistency.

### v3

@@ 4 Consistent conjectures
In this section, we present a sufficient condition on the family of conjectures, $\Phi$, such that the set of $\Phi$-solutions is nonempty (Theorem 4.1). We dub the conjectures satisfying the sufficient condition as consistent conjectures. Consistent conjectures guarantee candidate matchings, defined below, are a solution. In what follows, we first introduce the candidate matchings, and then we state Theorem 4.1. After discussing the intuition behind the result, we connect it to similar observations in Hafalir (2008) and Doval (2022). We conclude this section by showing the extension to dynamic matching markets of a standard solution concept in matching with externalities, rational expectations, does not entail consistent conjectures.
@@
    Utilities $\{u_{\varphi_{1}}(a,\cdot):B_{1}\cup\{a\}\rightarrow\mathbb{R}|a\in A_{1}\},\{v_{\varphi_{1}}(\cdot,b):A_{1}\cup\{b\}\rightarrow\mathbb{R}|b\in B_{1}\}$ are given by:
@@
      For each agent $a\in A_{1}$, $u_{\varphi_{1}}(a,a)=\min\{U_{1}(a,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,a)\}$. Moreover, $u_{\varphi_{1}}(a,b)=u(a,b)$ for $b\in B_{1}$.
@@
      For each agent $b\in B_{1}$, $v_{\varphi_{1}}(b,b)=\min\{V_{1}(b,\overline{m}):\overline{m}\in\varphi_{1}(\emptyset,b)\}$. Moreover, $v_{\varphi_{1}}(a,b)=v(a,b)$ for $a\in A_{1}$.
@@
In the one-period economy induced by the conjectures $\varphi_{1}$, an agent $a$ ranks matching partners on side $B$ according to $u(a,\cdot)$, except that agents on side $B$ who are worse than the worst conjectured continuation matching are now deemed unacceptable. In other words, in the one-period economy induced by $\varphi_{1}$, each agent in $A_{1}\cup B_{1}$ has the worst conjectured matching as an outside option.
@@
Suppose we know $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence and consider the following set of candidate matchings:
@@
In words, a candidate matching, $m^{\star}\in M^{\star}$, satisfies the following. In period $1$, $m_{1}^{\star}$ is a stable matching for $E_{1,\varphi_{1}}$, that is, $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$. Because the one-period economy induced by $\varphi_{1}$ is devoid of any externalities, its set of stable matchings is nonempty, and hence, at least one such matching exists (Gale and Shapley, 1962). Given the period-1 matching, $m^{\star}_{1}$, the continuation matching, $(m^{\star}_{t})_{t=2}^{T}$ is selected from $\Sigma_{T-1}^{\Phi}(E_{2}^{T}(m^{\star}_{1}))$, which is nonempty by assumption. The set of candidate matchings may not be a singleton either because $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$ is not a singleton, or because, for a given stable matching in $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, multiple continuation matchings may be solutions in the induced economy.Footnote 8Footnote 8Footnote 8Whereas many stable matchings for the one-period economy induced by the period-1 conjectures may exist, the Lone Wolf Theorem (McVitie and Wilson, 1970) implies they all have the same set of unmatched agents, and hence, they all induce the same continuation economy.
Any matching in $M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$ is a natural candidate to be a $\Phi$-solution because it satisfies most of the conditions in 7. To illustrate, fix any such matching, $m^{\star}$, and note the following. First, $m^{\star}$ satisfies part 4 of 7 by construction. Second, $m^{\star}$ satisfies parts 1, 2, and 3 of 7 for all period-1 agents $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)\neq k$. By definition of $E_{1,\varphi_{1}}$, every agent who matches in $t=1$ is matched to someone better than their worst conjectured matching. Furthermore, because $m^{\star}_{1}$ is chosen from $\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, no pair of matched agents under $m^{\star}_{1}$ can form a blocking pair.
@@
As we explain next, the only reason any such matching $m^{\star}$ may fail to be a $\Phi$-solution is that a “gap” exists between the conjectures of an agent who is unmatched in $t=1$ under $m^{\star}$ and the matching $m^{\star}$. To illustrate, note the above discussion implies that if $m^{\star}$ is not a $\Phi$-solution, $m^{\star}$ has a block involving a period-$1$ agent, $k$, who is unmatched in $t=1$, that is, $m^{\star}_{1}(k)=k$. We claim next that if $k$ blocks $m^{\star}$, then $m^{\star}$ is not a conjecture for $k$. To fix ideas, let $k=a\in A_{1}$ and suppose $a$ blocks $m^{\star}_{1}$ with an agent $b\in B_{1}$. Because $m^{\star}_{1}\in\mathrm{S}\left(E_{1,\varphi_{1}}\right)$, it must be that at least one of $a$ or $b$ prefers their worst conjectured matching to matching with each other; assume for simplicity this holds for $a$. Because $a$ wants to block $m^{\star}$ with $b$, it must be that $a$’s outcome under $m^{\star}$ is worse than the worst matching in $\varphi_{1}(\emptyset,a)$. In other words, $m^{\star}$ is not an element of agent $a$’s conjectures. Because $a$ cares only about the continuation matching $(m^{\star}_{t})_{t=2}^{T}$, that $m^{\star}$ is not part of $a$’s conjectures is akin to agent $a$’s conjectures not being consistent with the solution concept in the continuation. If, instead, $a$’s conjectures were consistent with the solution concept in the continuation, they would not be willing to block $m^{\star}$ with $b$.
The above discussion motivates the following definition:
Definition 9 (Consistent conjectures).
Suppose $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence and fix an economy $E^{T}$. Conjectures $\{\varphi_{1}(\emptyset,k):k\in A_{1}\cup B_{1}\}$ are consistent if for all candidate matchings $m^{\star}\in M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})$, the following holds:
for all kA1B1 such that m1(k)=k, we have that $\text{for all $k\in A_{1}\cup B_{1}$ such that $m^{\star}_{1}(k)=k$, we have that }m^{\star}\in\varphi_{1}(\emptyset,k).$ (CC)
In words, conjectures are consistent if each candidate matching, $m^{\star}$, is common to the conjectures of all agents who are unmatched under $m^{\star}$ in period $1$. For a given candidate matching $m^{\star}$, property CC places a constraint on the conjectures of those agents who are unmatched in period $1$ under $m^{\star}$: all of them should agree $(m^{\star}_{t})_{t=2}^{T}$ is a valid continuation matching. Instead, property CC places no restriction on those agents who are matched in period $1$ under $m^{\star}$: indeed, these agents may have different conjectures about the matching that would ensue conditional on remaining unmatched.
We have the following result:
@@
Suppose $\Sigma_{T-1}^{\Phi}$ is a nonempty-valued correspondence. Fix $E^{T}$ and suppose the conjectures $\{\varphi_{1}(\emptyset,k):k\in A_{1}\cup B_{1}\}$ are consistent. Then, $\Sigma_{T}^{\Phi}(E^{T})$ is nonempty. In particular, $M^{\star}(\varphi_{1},\Sigma_{T-1}^{\Phi})\subset\Sigma_{T}^{\Phi}(E^{T})$.
In words, consistency of the conjectures guarantees each of the candidate matchings are solutions for the economy. The proof of Theorem 4.1, which can be found in Appendix A, shows a stronger result: if property CC holds for some candidate matching, $m^{\star}$, then $m^{\star}$ is a solution for the economy. That is, to show the set of $\Phi$-solutions for $E^{T}$ is nonempty, assuming property CC holds for some candidate matching is enough. Because the nonempty-valued solution concepts we introduce in this and the following section satisfy the stronger property in 9, we prefer to make that our definition of consistency.
@@
As stated, the condition in Theorem 4.1 seems difficult to check without actually computing the set of candidate matchings. Yet, Theorem 4.1 suggests conjectures that always lead to nonempty solutions without actually computing $m^{\star}$. Indeed, it follows from Theorem 4.1 that if each period-1 agent conjectures that from $t=2$ onward, the continuation matching is an element of $\Sigma_{T-1}^{\Phi}(\cdot)$, $\Sigma_{T}^{\Phi}$ is nonempty valued. Formally, define for each $k\in A_{1}\cup B_{1}$,
$\varphi_{1}^{*}(\emptyset,k)=\left\{\overline{m}\in\pazocal{M}_{T}(\emptyset,k):(\overline{m}_{t})_{t=2}^{T}|_{E_{2}^{T}(\overline{m}_{1})}\in\Sigma_{T-1}^{\Phi}(E_{2}^{T}(\overline{m}_{1}))\right\}.$ (4.1)
@@
These conjectures reflect agent $k$ believes the economy abides by the set of $\Phi$-solutions from period $2$ onward, akin to a perfection requirement. For instance, if $T=2$, Equation 4.1 states agent $k$ anticipates the continuation matching, $\overline{m}_{2}$, is a stable matching in the induced economy, $E_{2}^{2}(\overline{m}_{1})$. That these conjectures induce a nonempty solution concept has two interesting implications. First, considering larger conjectures (in the set-inclusion sense) than those in Equation 4.1 is not needed to guarantee existence. The latter is a desirable property because a larger set of conjectures than those in Equation 4.1 would include matchings whose continuations do not satisfy the solution concept and are, in a sense, non-credible. Second, it implies the perfection requirement alone is compatible with the existence of solutions for dynamic matching markets. Instead, difficulties may arise if the conjectures also place structure on the period-$t$ matching, as we do in Section 5.
@@
Even if the statement of Theorem 4.1 relies on the recursive definition of the set of $\Phi$-solutions for a length-$T$ economy taking the nonemptiness of $\Sigma_{T-1}^{\Phi}$ as given, the procedure to construct $m^{\star}$ and the statement of Theorem 4.1 can be expressed solely in terms of the model primitives. We do this in Appendix A for completeness. When expressed in terms of the model primitives, consistency requires property CC holds in each period along the path of $m^{\star}$. That is, consistency is defined as the requirement that all candidate matchings, $m^{\star}$, satisfy that for all $t\geq 1$, all $k\in\pazocal{A}_{t}(m^{\star^{t-1}})\cup\pazocal{B}_{t}(m^{\star^{t-1}})$ such that $m^{\star}_{t}(k)=k$, then $m^{\star}\in\varphi_{t}(m^{\star^{t-1}},k)$ (see 14). In the same appendix, we formally define the conjectures and the solution concept suggested by Equation 4.1. That is, we iteratively apply Equation 4.1, starting from length-1 economies, to define a solution concept with conjectures consisting of all matchings whose continuations are solutions in the induced continuation economy. By Theorem 4.1, this solution concept is always nonempty (see A.1 in Appendix A).Footnote 9Footnote 9Footnote 9The solution concept defined in Appendix A is the analogue for irreversible matching markets of the one introduced in Kotowski (2019) for repeated matching markets.
@@
Connection with Hafalir (2008) and Doval (2022) Hafalir (2008) studies stability in two-sided, one-to-one, matching markets with payoff externalities. The starting point of his analysis is that the negative results in Sasaki and Toda (1996) follow from conjectures being exogenously defined without regard to the economy at hand. The first main result in Hafalir (2008) provides a sufficient condition on the conjectures for a stable matching to exist in his environment. To explain the condition, we review the construction in Hafalir (2008). Using the conjectures, Hafalir (2008) constructs a two-sided economy without externalities as in Sasaki and Toda (1996). In present notation, consider a one-period economy with agents in $A_{1}\cup B_{1}$. For each pair of agents, $(a,b)\in A_{1}\times B_{1}$, a conjecture $\varphi(a,b)$ is a subset of the set of matchings such that $(a,b)$ is matched.Footnote 10Footnote 10Footnote 10Hafalir (2008) does not consider blocks by individual agents. Given the conjectures, say $a$ prefers $b$ to $b^{\prime}$ if $a$ prefers the worst matching in $\varphi$$(a,b)$ to the worst matching in $\varphi(a,b^{\prime})$. The preferences for agents on side $B$ are similarly defined. Under these induced preferences, a stable matching is guaranteed to exist. Fix any such matching, $\overline{m}$. Hafalir (2008, Proposition 2) shows that if for all pairs $(a,b)$ such that $\overline{m}(a)=b$, we have that $\overline{m}\in\varphi(a,b)$, then $\overline{m}$ is stable in the economy with externalities.
@@
Theorem 4.1 identifies a key property that underlies the existence result in Doval (2022). In fact, Remark 2 in that paper points out that the conjectures corresponding to dynamic stability satisfy condition CC for any dynamically stable matching—not just the corresponding candidate matchings $m^{\star}$—and discusses the key role this property plays in the existence result. In Doval (2022), the conjectures are defined recursively based on what the solution concept specifies in the continuation, similar to the conjectures in Equation 4.1. Because our framework isolates the conjectures from the solution concept, it allows us to identify the role of consistency without reference as to how the conjectures relate to the solution concept. In doing so, the result offers an alternative way to construct solution concepts for dynamic matching markets by focusing on guaranteeing the conjectures are consistent. Section 5 illustrates this approach.
@@
First, considering the implications of not having consistent conjectures in the context of the candidate matchings is instructive. To fix ideas, consider one such $m^{\star}$ and suppose an unmatched agent on side $A$, $a$, wishes to block $m^{\star}$ with an agent on side $B$, $b$. Suppose also that $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$, and prefers matching with agent $b$ to $m^{\star}$, so that $m^{\star}\notin\varphi_{1}(\emptyset,a)$. Agent $a$’s block is in a sense not credible: whereas $a$ prefers matching with $b$ to $m^{\star}$, as soon as $a$ matches with $b$, $a$ would prefer to break up with $b$. After all, $a$ prefers the worst matching in $\varphi_{1}(\emptyset,a)$ to matching with $b$. Thus, whenever the failure of consistency rules out a candidate matching $m^{\star}$ as a solution, this may be because the solution concept allows for these non-credible blocks.
To make the second point, considering a more general definition of consistency that dispenses with the candidate matchings is useful. Given a family of conjectures, $\Phi$, and the induced solution concept $\{\Sigma_{t}^{\Phi}:t\in\{1,\dots,T\}\}$, we say the solution concept satisfies generalized consistency if the following holds. For all economies $E^{T}$ and $\Phi$-solutions $m$ for $E^{T}$, for all periods $t$, and $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$, if $m_{t}(k)=k$, then $m\in\varphi_{t}(m^{t-1},k)$. That is, we require that condition CC holds not necessarily for the candidate matchings, which may or may not be a solution, but for the actual solutions (provided they exist).
Suppose we know the set of $\Phi$-solutions is nonempty, but does not satisfy generalized consistency. For concreteness, let $m$ be a solution for $E^{T}$ for which a period $t\geq 1$ and an agent $k\in\pazocal{A}_{t}(m^{t-1})\cup\pazocal{B}_{t}(m^{t-1})$ exist such that $m_{t}(k)\neq k$ and $m\notin\varphi_{t}(m^{t-1},k)$. Because $m$ is a solution, the issue of credible blocks discussed above does not arise. However, that $k$ does not conceive of $m$ as a potential outcome for the continuation economy suggests a tension exists between what the solution concept prescribes in the continuation economy in which $k$ is one of the agents available to match, and what $k$ expects would happen if they would object to $m$. That $m$ is not part of $k$’s conjectures means $k$ believes $m$ cannot be enforced in the continuation. Thus, if the solution concept captures the set of self-enforcing outcomes in the continuation economy and $k$ agrees with this assessment, $m$ should be part of $k$’s conjectures and the solution concept should satisfy generalized consistency. As we illustrate below, the rational expectations solution fails both consistency and generalized consistency.